# Competing Mechanisms at Vibrated Interfaces of Density-Stratified Fluids

### v1

@@ Abstract
Fluid–fluid interfacial instability and subsequent fluid mixing are ubiquitous in nature and engineering. Two hydrodynamic instabilities have long been thought to govern the interface behavior: the pressure gradient-driven long-wavelength Rayleigh–Taylor (RT) instability and resonance-induced short-wavelength Faraday instability. However, neither instability alone can explain the dynamics when both mechanisms act concurrently. Instead, we identify a previously unseen multi-modal instability emerging from their coexistence. We show how vibrations govern transitions between the RT and Faraday instabilities, with the two competing instead of resonantly enhancing each other. The initial transient growth is captured by the exponential modal growth of the most unstable Floquet exponent, along with its accompanying periodic behavior. Direct numerical simulations validate these findings and track interface breakup into the multiscale and nonlinear regimes, in which the growing RT modes are shown to suppress Faraday responses via a nonlinear mechanism.
@@
The interface separating phases in multiphase fluid systems is often subjected to deformation due to internal density differences or external vibrations. These deformations manifest in various forms, including liquid dripping [1, 2, 3], bubble injection [4, 5, 6], tip streaming [7, 8, 3], and surface waves [9]. Such phenomena are ubiquitous, occurring in natural processes such as rainfall and tidal movements, as well as in engineering applications like atomization and fusion reactions [10, 11]. The unstable growth of these interface deformations eventually leads to interface breakup and fluid mixing. Linear stability analysis predicts two primary hydrodynamic instability mechanisms: the pressure gradient-driven fluid mixing mechanism, triggering the Rayleigh–Taylor (RT) instability [12, 13], and the resonance mechanism, triggering a Faraday instability [14]. Over a century of research has focused on studying these fundamental hydrodynamic instabilities separately, as each is central to a vast range of applications (for RT see [15, 16, 17, 18, 19, 20] and for Faraday see [21, 9, 22, 23, 24, 25, 26]).
@@
Despite their well-established nature and pervasive observations, a comprehensive understanding of the interface’s behavior when the long-wavelength RT and short-wavelength Faraday instabilities coexist is lacking. This gap is crucial, as vibrations and density stratification often occur simultaneously in realistic situations. Wolf [27, 28] showed via experiment that unstable viscous RT waves can be stabilized without generating standing Faraday waves by oscillating the container at specific amplitudes and frequencies. Later studies supported this conclusion through theoretical arguments, and a partial theory was developed [29, 30, 31, 32, 33]. However, the coexistence of unstable RT and Faraday modes, where they actively influence the interface dynamics, has been largely overlooked. The complex interactions and physical mechanisms emerging from this coexistence remain mysterious. This work addresses this gap and provides a comprehensive view of interfacial dynamics under combined density stratification and vibrations.
@@
We analyze the coexistence and competition of Rayleigh-Taylor (RT) and Faraday instabilities under the effect of external vibrations and internal density differences. We examine the temporal evolution of the interface within the linear regime. Using horizontal homogeneity, we decompose the interface into individual wave numbers, allowing for the investigation of both long-wavelength RT and short-wavelength Faraday waves. Two-dimensional scale-resolved numerical simulations validate the linear theory and capture the system’s evolution into multiscale and nonlinear regimes. We also explore the oscillation–frequency-amplitude phase space to characterize the transition of the dominant instability.
@@
$\tilde{g}(t)=g^{*}g+a\cos(\omega t+\varphi_{0}),$ (1)
where $a$ is the oscillatory amplitude, $\omega$ is the corresponding frequency, and $\varphi_{0}$ is the initial phase. Positive gravity, $g^{*}=1$, applies when the denser fluid is at the domain bottom, and negative gravity, $g^{*}=-1$, applies when the denser fluid is on top. Figure 1 shows the two-fluid interface subjected to a harmonic vibration. The time-periodic variation in gravity is suspected to be the driving source of instability, motivating Floquet analysis [34]. We define nondimensionalized quantities, denoted by $(\cdot)^{*}$, in table A.1 in the End Matter. The temporal evolution of the interface displacement, $\zeta^{*}(t)$, is decomposed into an exponential term, $\mathrm{e}^{\gamma^{*}t^{*}}$, that models the growth or decay and a periodic term, $\sum_{n}\hat{\zeta}_{n}^{*}\mathrm{e}^{\mathrm{i}n\omega^{*}t^{*}}$, which accounts for the oscillatory fluctuations about it.
We assume that the interface profile is well-defined and single-valued. At each horizontal wavenumber, ${k^{*}}$, a separation into amplitude growth (or decay) and periodic oscillations yields
@@
$\hat{\zeta}^{*}(k^{*},t^{*})=\underbrace{\mathrm{e}^{\gamma^{*}t^{*}}}_{\text{Modal}}\,\underbrace{\sum_{n=-\infty}^{\infty}\hat{\zeta}_{n}^{*}(k^{*})\mathrm{e}^{\mathrm{i}n\omega^{*}(t^{*}+t_{0}^{*})}}_{\text{Periodic components}},$ (2)
where $\gamma^{*}$ is the Floquet exponent, $t_{0}^{*}$ is the time shift associated with the initial phase $\varphi_{0}$, and $n$ is the index of the harmonics [23, 24]. A positive real Floquet exponent indicates instability, driving the system toward interface breakup. Floquet exponents, often obtained through the eigenvalue decomposition of the monodromy matrix, are not unique and collectively describe nonmodal growth. Here, we demonstrate that the modal growth of the most unstable Floquet exponent, $\gamma^{*}_{U}$, combined with the periodic components, accurately predicts the initial transient dynamics. By relating the interface displacements at different orders of harmonics, $\gamma^{*}_{U}$ is found for each wavenumber by solving
$\gamma^{*}_{U}=\operatorname*{arg\!\max}_{\det{{\bf\it A}(\gamma^{*};\,k^{*})-a^{*}{\bf\it B}}=0}\mathrm{Re}\{\gamma^{*}\}.$ (3)
@@
The periodic components at different harmonics, $\hat{{\bf\it\zeta}}^{*}$, can be then found as the null space of $({\bf\it A}-a^{*}{\bf\it B})$. A detailed analysis is provided in the End Matter. Equation 3 shows a generalized analysis of the hydrodynamic instability in fluid mixing, covering a wide range of configuration parameters, including viscosity, surface tension, vibration, and density differences. In the static limit ($a^{*}\to 0$), Eq. 3 becomes
@@
for the Faraday instability [23, 24]. Beyond the identification of the stability boundary, Eq. 2 enables the prediction of transient wave dynamics at a given wavenumber $k^{*}$ within the unstable regime. Together, these analyses enable the dissection of the mechanics as the system transitions toward nonlinear interface breakup. An example of this utility is shown next.
IIIResults

### v2

# Competing Mechanisms at Vibrated Interfaces of Density-Contrast Fluids

@@ Abstract
Fluid–fluid interfacial instability and subsequent fluid mixing are ubiquitous in nature and engineering. The hydrodynamic instability of fluid interfaces has long centered on the pressure gradient-driven long-wavelength Rayleigh–Taylor instability and the resonance-induced short-wavelength Faraday instability. However, neither instability alone can explain the dynamics when both mechanisms are present. We identify a previously unseen multi-modal instability emerging from their coexistence. When the denser fluid is polydimethylsiloxane, the mixed region at a high density contrast (Atwood number=0.9) spans a vibration amplitude range approximately twice the gravitational acceleration. Using Floquet stability analysis, we show how vibrations govern transitions between the RT and Faraday instabilities, leading to contention between these instabilities rather than resonant enhancement. The initial transient growth is represented by the exponential modal growth of the most unstable Floquet exponent, along with its accompanying periodic behavior. Direct numerical simulations validate these findings and track interface breakup into the multiscale and nonlinear regimes. Specifically, we show that growing RT modes nonlinearly suppress Faraday responses even when the initial growth rate of the Faraday instability is 3.63 times that of RT, so a bidirectional competition hinders their sustained coexistence.
@@ I Introduction
The interface separating phases in multiphase fluid systems is often subjected to deformation due to internal density differences or external vibrations. These deformations manifest in various forms, including liquid dripping [1, 2, 3], bubble injection [4, 5, 6], tip streaming [7, 8, 3], and surface waves [9]. Such phenomena are ubiquitous, occurring in natural processes such as mantle plumes and cryospheric vibrations, as well as in engineering applications like atomization and inertial confinement fusion [10, 11]. The unstable growth of these interface deformations eventually leads to interface breakup and fluid mixing. Linear stability analysis predicts two primary hydrodynamic instability mechanisms: the pressure gradient-driven fluid mixing mechanism, triggering the Rayleigh–Taylor (RT) instability [12, 13], and the resonance mechanism, triggering a Faraday instability [14]. Over a century of research has focused on studying these fundamental hydrodynamic instabilities separately, as each is central to a vast range of applications.
The analysis of RT instability originated with investigations into the growth of small-amplitude interfacial waves between two inviscid fluids of differing densities [12, 13]. Viscous effects in the stable configuration were first considered by Harrison [15], and later, Bellman and Pennington [16] examined the combined influence of viscosity and surface tension on the unstable regime. A comprehensive and widely acclaimed treatise on RT instability can be found in Chandrasekhar [17]. Building on this foundation, Plesset and Whipple [18] introduced simplified physical models applicable in the asymptotic limits of large and small contrasts in density and viscosity. For detailed reviews of RT instability theory and its developments, readers are referred to Sharp [19] and Kull [20]. More recently, Piriz et al. [21] derived a simple yet highly accurate analytical expression for RT instability in nonideal fluids.
When vibration amplitude exceeds a critical threshold, the flat interface destabilizes through resonant Faraday instability, generating standing surface waves. This phenomenon has garnered significant attention and extensive study; see an early review by Miles and Henderson [9]. Benjamin and Ursell [22] showed that the stability of the free surface of an ideal fluid can be analyzed using a system of Mathieu equations. Although widely adopted, this idealized model is not sufficiently realistic to yield results that match experimental observations [23]. This discrepancy shows the role of viscosity in interfacial stability. Using Floquet analysis, Kumar and Tuckerman [24] showed that even when linear damping is included, the Mathieu equation fails to represent the dynamics of viscous fluids. They demonstrated that viscosous effects modify the wavelength selection and distort the tongue-shaped stability regions. Kumar [25] extended this analysis to shallow fluids with free surfaces, predicting a series of bicritical points for subharmonic and harmonic responses. Despite its limitations in representing viscous effects, Wright et al. [26] numerically demonstrated that the modified Mathieu equation accurately predicts the temporal growth of Faraday oscillations in inviscid fluids. Beyond the waveforms observed in the above two-dimensional analyses and experiments, Panda et al. [27] recently demonstrated numerically that Faraday instability can give rise to more complex patterns when the domain is extended to three dimensions.
@@
Despite their well-established nature and pervasive observations, a comprehensive understanding of the interface behavior when the long-wavelength RT and short-wavelength Faraday instabilities coexist is lacking. This knowledge gap is important because vibrations and density contrast occur simultaneously in many realistic situations, including vibrated multiphase systems, industrial mixing processes, and fluids in spacecraft propellant tanks. For example, vibrated levitating liquids represent a compelling case in which these instabilities interact in a nontrivial way [28]. Faraday waves inherently link to RT instability when the oscillation amplitude exceeds gravitational acceleration [29, 26]. In such regimes, the interface undergoes the Faraday instability throughout each oscillation cycle, while experiencing RT mechanisms during the upward acceleration phase. Notably, Dinesh et al. [30] experimentally demonstrated the emergence of an electrostatic resonance resembling an RT-like instability, coexisting with the standard Faraday instability. A fundamental understanding of these coupled mechanisms is essential for developing effective strategies to control interfacial dynamics in such complex systems. Wolf [31, 32] showed via experiment that unstable viscous RT waves can be stabilized without generating standing Faraday waves by oscillating the container at specific amplitudes and frequencies. Later studies supported this conclusion through theoretical arguments, and a partial theory was developed [33, 34, 35, 36, 37]. However, the coexistence of unstable RT and Faraday modes, which can meaningfully influence the interface dynamics, has been largely overlooked. The complex interactions and physical mechanisms emerging from this coexistence remain mysterious. This work addresses this gap and provides a comprehensive view of interfacial dynamics under combined density contrast and vibrations, while also accounting for viscosity and surface tension effects.
@@
We focus on two-dimensional analyses and simulations to gain clearer insight into the underlying mechanisms. Figure 1 shows a schematic of the two-fluid interface, where both long-wavelength RT and short-wavelength Faraday waves coexist. In section II, we perform a linear Floquet stability analysis to characterize the temporal evolution of the interface in the linear regime. Assuming horizontal homogeneity, we decompose the interface into individual wave numbers, enabling separate investigation of both long-wavelength RT and short-wavelength Faraday modes. Section III describes the numerical approach for two-dimensional, scale-resolved simulations of the two-phase fluid system. These simulations validate the linear theory and represent the system evolution into multiscale and nonlinear regimes, as presented in section IV. We also explore the oscillation–frequency-amplitude phase space to characterize the transition of the dominant instability. Section V discusses the limitations of the study and summarizes the findings on the coexistence and competition between RT and Faraday instabilities.
@@ II Theoretical Background
II.1 Governing equations
Consider the interface between two immiscible and incompressible fluids: a denser and lighter fluid, denoted by subscripts $\left(\cdot\right)_{d}$ and $\left(\cdot\right)_{l}$. The fluids are confined between two infinite horizontal plates separated by a distance of $2h$ and subject to constant gravitational acceleration and an oscillatory vertical acceleration
$\tilde{g}(t)=g^{*}g+a\cos(\omega t+\varphi_{0}),$ (1)
where $a$ is the oscillatory amplitude, $\omega$ is the corresponding frequency, and $\varphi_{0}$ is the initial phase. Positive gravity, $g^{*}=1$, applies when the denser fluid is at the domain bottom, and negative gravity, $g^{*}=-1$, applies when the denser fluid is on top. Figure 1 shows the two-fluid interface subjected to a harmonic vibration. Within each fluid layer $j$, the motion of the fluid is governed by the incompressible Navier–Stokes equations,
$\rho_{j}\left[\partial_{t}+\left({\bf\it u}_{j}\cdot{\bf\it\nabla}\right)\right]{\bf\it u}_{j}=-\gradient p_{j}-\rho_{j}\tilde{g}{\bf\it e}_{z}+\mu_{j}\gradient^{2}{\bf\it u}_{j},\quad\nabla\cdot{\bf\it u}_{j}=0.$ (2)
To systematically analyze the interfacial dynamics, we nondimensionalize the governing equations using $h$ as the characteristic length and $g$ as the reference gravitational acceleration. Density, $\rho$, viscosity, $\mu$, and surface tension, $\sigma$ are scaled using appropriate reference values. The resulting dimensionless variables, denoted by $(\cdot)^{*}$, are summarized in table 1.
Table 1: Dimensionless quantities.
Dimensionless quantity Quantity name
$t_{c}=\sqrt{h/g}$ Gravitational time
$\mathrm{At}=(\rho_{d}-\rho_{l})/(\rho_{d}+\rho_{l})$ Atwood number
$\eta=\mu_{l}/\mu_{d}$ Viscosity ratio
$C=\nu_{d}/\sqrt{gh^{3}}$ Viscous-to-gravitational
force ratio
$\mathrm{Bd}=\rho_{d}gh^{2}/\sigma_{d}$ Bond number
$a^{*}=a/g$ Acceleration
$\zeta^{*}=\zeta/h$ Interface displacement
$k^{*}=kh$ Wavenumber
$t^{*}=t/t_{c}$ Time
$\gamma^{*}=\gamma t_{c}$ Growth rate
$\omega^{*}=\omega t_{c}$ Oscillatory frequency
Eq. 2 can therefore be written as
$\left[\partial_{t^{*}}+\left({\bf\it u}^{*}_{j}\cdot{\bf\it\nabla}^{*}\right)\right]{\bf\it u}^{*}_{j}$ $=-\gradient^{*}p^{*}_{j}-\tilde{g}^{*}{\bf\it e}_{z}+C_{j}{\bm{\Delta}^{*}}{\bf\it u}^{*}_{j},\quad\gradient^{*}\cdot{\bf\it u}_{j}^{*}=0,$ (3)
where $C_{d}=C$, and $C_{l}=(1-\mathrm{At})/(\eta(1+\mathrm{At}))$. We can decompose the flow state around the equilibrium-state solution of the Navier–Stokes equations 3 as
${\bf\it u}^{*}_{j}={\bf\it U}^{*}_{j}+{{\bf\it u}^{*}_{j}}^{\prime},\quad p^{*}_{j}={P}^{*}_{j}+{p^{*}_{j}}^{\prime},$ (4)
where $({\bf\it U}^{*}_{j},\,P^{*}_{j})=(0,-\tilde{g}^{*}{\bf\it e}_{z})$ and $(\cdot)^{\prime}$ denotes the small fluctuating components. The linearized governing equations for the fluctuations are
$\partial_{t^{*}}{{\bf\it u}_{j}^{*}}^{\prime}=-\gradient^{*}{p^{*}_{j}}^{\prime}+C_{j}{\bm{\Delta}^{*}}{{\bf\it u}_{j}^{*}}^{\prime},\quad\gradient^{*}\cdot{{\bf\it u}_{j}^{*}}^{\prime}=0.$ (5)
Applying the operator ${\bf\it e}_{z}\cdot\gradient^{*}\times\gradient^{*}\times$ to Eq. 5 eliminates the horizontal velocity components, yielding
$\left(\partial_{t^{*}}-C_{j}{\bm{\Delta}^{*}}\right){\bm{\Delta}^{*}}{w_{j}^{*}}^{\prime}=0,$ (6)
At the interface $z^{*}=\zeta^{*}$, continuity of the vertical velocity perturbation, $w_{j}^{*}$$^{\prime}$, and the tangential stress yields
$\delta\left[{w_{j}^{*}}^{\prime},\,\partial_{z^{*}}\left({w_{j}^{*}}^{\prime}\right),\,C_{j}\left(\bm{\Delta}^{*}-2\partial_{z^{*}z^{*}}\right){w_{j}^{*}}^{\prime}\right]=0,$ (7)
where $\delta$ represents the jump across the interface. The linearized kinematic boundary condition gives
$\partial_{t^{*}}{\zeta}^{*}$ $={w}^{*}|_{z^{*}=0}.$ (8)
The jump condition for pressure across the interface,
$\delta p^{*}$ $=2(C_{l}-C_{d})\left(\partial_{z^{*}}{w}^{*}\right)|_{z^{*}=0}-\frac{2\mathrm{At}}{1+\mathrm{At}}\tilde{g}^{*}\zeta^{*}+\frac{1}{\mathrm{Bd}}\left(\bm{\Delta}^{*}-\partial_{z^{*}z^{*}}\right)\zeta^{*},$ (9)
explicitly represents the effect of vibration. Together, Eqs. 6 to 9 govern the linearized interfacial dynamics. We further assume that the interface profile, $\zeta^{*}$, remains well-defined and single-valued during the vibration. Due to the horizontal homogeneity, we use a Fourier transform in the horizontal direction,
@@ II.1 Governing equations
to expand the velocity and interface displacement in the above governing equations. Eq. 10 decomposes waveforms into different length scales, enabling separate analysis of RT and Faraday instabilities. The dynamics introduced by the parametric source are subsequently analyzed using a Floquet analysis, as described next.
II.2 Linear Floquet stability analysis
The time-periodic variation in gravity is suspected to be the driving source of instability, thereby motivating a Floquet analysis [38]. Consequently, the temporal evolution of the vertical interface displacement, $\zeta^{*}(t^{*})$, is expected to exhibit both exponential growth or decay, modeled by $\mathrm{e}^{\gamma^{*}t^{*}}$, and superimposed oscillatory fluctuations, captured by the periodic term $\sum_{n}\hat{\zeta}^{*}_{n}\mathrm{e}^{\mathrm{i}n\omega^{*}t^{*}}$. At each horizontal wavenumber, this response can be decomposed into modal and periodic parts, yielding
@@ II.2 Linear Floquet stability analysis
$\left[\hat{w}^{*}_{j}(k^{*},z^{*},t^{*}),\,\hat{\zeta}^{*}(k^{*},t^{*})\right]=\underbrace{\mathrm{e}^{\gamma^{*}t^{*}}}_{\text{Modal}}\,\underbrace{\sum_{n=-\infty}^{\infty}\left[\hat{w}_{j,n}^{*}(k^{*},z^{*}),\,\hat{\zeta}_{n}^{*}(k^{*})\right]\mathrm{e}^{\mathrm{i}n\omega^{*}(t^{*}+t_{0}^{*})}}_{\text{Periodic components}},$ (11)
where $\gamma^{*}$ is the Floquet exponent, $t_{0}^{*}$ is the time shift associated with the initial phase $\varphi_{0}$, and $n$ is the intger index of the harmonics [24, 25]. A positive real Floquet exponent indicates instability, driving the system toward interface breakup. Substituting Eqs. 10 and 11 into Eq. 6 gives
$\left[\gamma_{n}^{*}-C_{j}D_{zz}^{(k)}\right]D_{zz}^{(k)}\hat{w}^{*}_{n,j}=0,$ (12)
where $\gamma_{n}^{*}\equiv\gamma^{*}+\mathrm{i}n\omega^{*}$ is the total Floquet exponent for the $n$th harmonics, and $D_{zz}^{(k)}\equiv\partial_{z^{*}z^{*}}-{k^{*}}^{2}$. The general solution of this dispersion relation is
$\hat{w}^{*}_{n,j}$ $=a_{n,j}\mathrm{e}^{k^{*}z^{*}}+b_{n,j}\mathrm{e}^{-k^{*}z^{*}}+c_{n,j}\mathrm{e}^{k^{*}q_{n,j}z^{*}}+d_{n,j}\mathrm{e}^{-k^{*}q_{n,j}z^{*}},$ (13)
where $q_{n,j}\equiv\sqrt{1+\gamma_{n}^{*}/(C_{j}k^{*})}$. No-slip boundary conditions at the two plates yield
$\hat{w}^{*}_{n,d}=\partial_{z^{*}}\hat{w}^{*}_{n,d}=0\quad$ $\text{at}\quad z^{*}=-1,$ (14)
$\hat{w}^{*}_{n,l}=\partial_{z^{*}}\hat{w}^{*}_{n,l}=0\quad$ $\text{at}\quad z^{*}=1.$ (15)
The jump conditions at the interface in Eqs. 7 and 8 can be written as
$\left[1,\,\partial_{z^{*}},\,\eta\left(\partial_{z^{*}z^{*}}+{k^{*}}^{2}\right)\right]\hat{w}^{*}_{n,l}$ $=\left[1,\,\partial_{z^{*}},\,\left(\partial_{z^{*}z^{*}}+{k^{*}}^{2}\right)\right]\hat{w}^{*}_{n,d},$ (16)
$\gamma_{n}^{*}\hat{\zeta}^{*}_{n}$ $=\hat{w}^{*}_{n}|_{z^{*}=0}.$ (17)
Substituting the general solution in Eq. 13 into the above 8-equation boundary condition, we find
${{\bf\it Q}_{n}}\matrixquantity(a_{n,d},b_{n,d},c_{n,d},d_{n,d},a_{n,l},b_{n,l},c_{n,l},d_{n,l})^{\mathsf{T}}=\matrixquantity(\gamma^{*}_{n}\hat{\zeta}^{*}_{n},0,0,0,0,0,0,0)^{\mathsf{T}}.$ (18)
The full expression of the matrix ${\bf\it Q}_{n}$ is given in Eq. Appendix in the appendix. When the depth $h\to\infty$, the vanishing velocity gives $b_{n,d}=d_{n,d}=a_{n,l}=c_{n,l}=0$.
The jump condition in Eq. 9 directly relates the displacements at different orders of harmonics, $\hat{\zeta}_{n}^{*}$. The resulting linear system can be expressed as
Eqs. LABEL:eqn:General_EVD and 19 provide a comprehensive analysis of interfacial instabilities, accounting for viscosity effects, surface tension, internal density differences, and external vibrations. For each wavenumber, the solution of the optimization problem
[…]
@@
yields the corresponding most unstable Floquet exponent. The periodic components at different harmonics, $\hat{{\bf\it\zeta}}^{*}$, can be then found as the null space of $({\bf\it A}-a^{*}{\bf\it B})$. Equation 20 shows a generalized analysis of the hydrodynamic instability in fluid mixing, covering a range of configuration parameters, including viscosity, surface tension, vibration, and density differences, exhibiting rich mechanical differences. In the static limit ($a^{*}\to 0$), Eq. 20 becomes
@@
for the Faraday instability [24, 25]. Beyond identifying the stability boundary, Eq. 11 enables the prediction of transient wave dynamics at a given wavenumber $k^{*}$ within the unstable regime. In practice, Eq. LABEL:eqn:General_EVD is truncated at $n=10$, resulting in a $22\times 22$ linear problem. Including higher-order harmonics does not visibly affect the results but may introduce numerical instabilities. Together, these analyses enable the dissection of the mechanics as the system transitions toward nonlinear interface breakup. While standard Floquet exponents, often obtained through the eigendecomposition of the monodromy matrix, are not unique, they collectively describe nonmodal growth. Using numerical simulations, we demonstrate that the modal growth of the most unstable Floquet exponent, $\gamma^{*}_{U}$, combined with the periodic components, can accurately predict the initial transient dynamics.
III Numerical method